OpenStudy (anonymous):

A roadside vegetable stand sells pumpkins for $5 each and squashes for $3 each. One day they sold 6 more squash than pumpkins, and their sales totaled $98. Write and solve a system of equations to find how many pumpkins and squash they sold? How do I do this?

OpenStudy (anonymous):

p = s - 6 5p + 3s = 98 now sub s - 6 in for p in the other equation and solve for s. Once s is found, sub that value back into either of the original equations and find x

OpenStudy (anonymous):

or you could use this as the first equation...s = p + 6. either one will work

OpenStudy (anonymous):

I'm confused.

OpenStudy (anonymous):

6 squashes cost $18 98-18=$80 80/(5+3)=10 pumpkins

OpenStudy (anonymous):

p = s - 6 now lets sub s - 6 in for p in the other equation 5p + 3s = 98 5(s - 6) + 3s = 98 5s - 30 + 3s = 98 8s - 30 = 98 8s = 98 + 30 8s = 128 s = 128/8 s = 16 now sub 16 in for s in either of the original equations p = s - 6 p = 16 - 6 p = 10 10 pumpkins and 16 squash Was that what you were having trouble with or was it just setting up the equations ?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."